OpenStudy (jamaljunior):

Amy and her sister are making lemonade for the school dance. Their recipe makes 2 quarts of lemonade. If they make 35 times the recipe, how many gallons of lemonade will Amy and her sister have?

563blackghost (563blackghost):

So conversion equations...got it! :) We know you can make `2 quarts of lemonade. Now they want to make that `35 times` so we would multiply 2 by 35....we will find the total quarts from that :) \(\huge{35 \times 2 = ?}\)

563blackghost (563blackghost):

Dont know if the LaTex is working for you so Ill retype in regular format.... So conversion equations...got it! :) We know you can make `2 quarts of lemonade`. Now they want to make that `35 times` so we would `multiply 2 by 35`....we will find the total quarts from that :) `35×2=?`

OpenStudy (jamaljunior):

70

563blackghost (563blackghost):

Correct :) Now we know it makes `70 quarts` but we need to know gallons...So how many quarts are in a gallon? (u should be able to access your conversion sheet) If not there is an attachment at the bottom :)

OpenStudy (jamaljunior):

280

OpenStudy (jamaljunior):

?

563blackghost (563blackghost):

So you are right about there being 4 quarts in a gallon since you multiplied but we wouldnt do that.... We would actually divide by 4 to find how many gallons.... \(\huge{\frac{70}{4}=?~gallons}\)

OpenStudy (jamaljunior):

ohhh 17.5

563blackghost (563blackghost):

\(\Huge{\checkmark}\)

563blackghost (563blackghost):

Correct :)

OpenStudy (jamaljunior):

ty

563blackghost (563blackghost):

Np :) if you ever see me on and need help just tag me and Ill get to you as soon as I can ^^